A group of men can do a job in as many days as there are men in the group. If the number of men is increased by 5, the number of days to do the job will be decreased by 4. Find the number of men in the group.
Answer by ankor@dixie-net.com(22740)

A group of men can do a job in as many days as there are men in the group.
Find the number of men in the group.
:
let m = no. of men in the group, also the no. of days to complete the job
Therefore
m^2 = no. of man-days to do the job
:
"If the number of men is increased by 5, the number of days to do the job will be decreased by 4."
(m+5)*(m-4) = m^2
FOIL
m^2 - 4m + 5m - 20 = m^2
m^2 - m^2 + m - 20 = 0
m = 20 men in the original group
:
:
See if that checks out:
20^2 = 400 man-days to do the job
(20+5)(20-4) =
25 * 16 = 400 man-days